OpenStudy (anonymous):

How many x-intercepts does the graph of y = x2 have?

OpenStudy (anonymous):

I have no clue how to solve this... and the sad thing is.... it's probably really easy...

OpenStudy (amistre64):

what is the value of y when we want to find a point on the x axis?

OpenStudy (anonymous):

@amistre64 0

OpenStudy (amistre64):

then, when y=0, we find the values for x y=x^2 0=x^2 when?

OpenStudy (anonymous):

we find the sqrt of 0 and x^2?

OpenStudy (amistre64):

we could, but this is simple enough: x^2 means: x*x x*x = 0 when x=?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

I hate zeros... too many rules

OpenStudy (amistre64):

yes, and since there is only one value of x that makes this statement true; then y=x^2 has only 1 x intercept

OpenStudy (anonymous):

okay I think I get it now

OpenStudy (amistre64):

there are harder ones, but the concept is the same :)

OpenStudy (anonymous):

okay thanks!